n + 4) (+3 Determine whether the following series converges or diverges. In (n + 5) a) b) In(+2) (-1) c) -- +1
Answer #1

= (-) In (+5) In(n+2) ln (nr 5 In (+2) lan) = فى lim lan) = lim ln(n+5) form no ennta) n-D In (nts) dn = lim (L Hospital nolim n+5 no nt2 n-do =lim n(1+2) n (its - lim 1 + 2 الحمد-۱ 145 1 + 6 to lim lan) = 1²0 n-D .. By Divergence test The Series SLet bn = 5 ntl an lim lim br no pru lim 5 It lim an = 5= finite (*) ug nob and Diverges R bon E I n=1 n=1 By Comparison test